What is the force per meter of length of a straight wire carrying a 3.40A current when placed at an angle of 40° to a 0.75T unifotm magnetic field?....(please help me by showing the working)
1
Expert's answer
2011-06-21T08:24:05-0400
Sinse F = B I L sin α, force per meter is

F/L = BI sin α = 0.75 * 3.4 * sin 40 = 1.64 N/m.